8x^2-x-7
THANK YOU!!!

Answer by kiru_khandelwal(79)

in a quadratic equation
ax^2+bx+c = 0
x = (-b + sqrt(b^2-4ac))/2a and (-b - sqrt(b^2-4ac))/2a
so for the quadratic equation
8x^2-x-7
a=8, b=-1, c=-7
therefore, x = (-(-1) + sqrt((-1)^2 - 4*8*(-7)))/2*8 = (1+sqrt(225))/16 =(1+15)/16=1
and also
x = (-(-1) - sqrt((-1)^2 - 4*8*(-7)))/2*8 = (1-sqrt(225))/16 =(1-15)/16=-14/16 = -7/8
so x = 1, -7/8